OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

\[if f(x)=3/5x and g(x)= -3x/2, find f(x) + g(x)\]

OpenStudy (alexandervonhumboldt2):

3/5x-3x/2=6/10x-15x/10=-9x/10=-0.9x

OpenStudy (mertsj):